OpenStudy (anonymous):

System of elimination- how do i find each system of equation of 6c - 9d = 111 and 5c - 9d =103?

OpenStudy (anonymous):

As we have a term that is the same in each equation (-9d), we can just subtract one equation from the other: \(6c-9d=111\) \(5c-9d=103\) \((6c-9d)-(5c-9d)=111-103\) Can you simplify the last equation please?

OpenStudy (mathmale):

Actually, you don't "find each system." The two given equations comprise (make up) a system. there's only one system. You mean "solve the following system of linear equations by elimination." I agree with Tom (above) -9d appears in both equations, so leave the first equation as is and subtract the second equation from the first. Enclose the 2nd equation in parentheses and then put a - sign in front of the left parenthesis. Please show your work. doing this will elilminate d and leave you with the task of finding only c.